# Translationese-index: Using Likelihood Ratios for Graded and Generalizable Measurement of Translationese

### v1

@@ Abstract
In this paper, we propose the first quantitative measure for translationese—the translationese-index (T-index) for graded and generalizable measurement of translationese, computed from the likelihood ratios of two contrastively fine-tuned language models (LMs). We use a synthesized dataset and a dataset with translations in the wild to evaluate T-index’s generalizability in cross-domain settings and its validity against human judgments. Our results show that T-index is both robust and efficient. T-index scored by two 0.5B LMs fine-tuned on only 1-5k pairs of synthetic data can well capture translationese in the wild. We find that the relative differences in T-indices between translations can well predict pairwise translationese annotations obtained from human annotators; and the absolute values of T-indices correlate well with human ratings of degrees of translationese (Pearson’s $r=0.568$). Additionally, the correlation between T-index and existing machine translation (MT) quality estimation (QE) metrics such as BLEU and COMET is low, suggesting that T-index is not covered by these metrics and can serve as a complementary metric in MT QE.
@@
Yikang Liu1thanks: Work done during an internship at Tongyi Lab. #Corresponding author, Wanyang Zhang3, Yiming Wang2, Jialong Tang4, Pei Zhang4, Baosong Yang4, Fei Huang4, Rui Wang2, Hai Hu1# 1School of Foreign Languages, Shanghai Jiao Tong University 2Dept. of Computer Science and Engineering, Shanghai Jiao Tong University 3School of Software and Microelectronics, Peking University 4Tongyi Lab Correspondence: yikangliu@sjtu.edu.cn; hu.hai@outlook.com
@@ 1 Introduction
Translationese refers to linguistic properties introduced in the translation process that are different from those of texts originally written in that language (Gellerstam, 1986). While such properties are not inherently undesirable, they often lead to unnatural and non-native-like language that differs from idiomatic and authentic texts.
Many attempts have been made to study translationese in the past decades. One approach trains binary classifiers between translated and original texts to uncover the features of translationese (Baroni and Bernardini, 2005; Koppel and Ordan, 2011; Volansky et al., 2015; Hu and Kübler, 2021; Pylypenko et al., 2021); another relies on the use of the distributional statistics of the features concerning translationese to evaluate the generated texts (Freitag et al., 2022; Guo et al., 2024; Li et al., 2025; Flamich et al., 2025).
Despite these efforts, so far, we do not yet have a quantitative measure for translationese to choose the more authentic and natural translation, out of many equally accurate translations, which is especially important when existing MT QE metrics focus more on the accuracy and become less reliable in evaluating MTs of higher quality (Agrawal et al., 2024; Kocmi et al., 2024). Quantifying the stylistic features at the sample level will facilitate more holistic MT QE.
Furthermore, there is a need to isolate translationese from genre-specific features of a text. Specifically, previous studies on translationese are often situated in a binary text-classification paradigm between translated and non-translated text (Baroni and Bernardini, 2005; Volansky et al., 2015; Pylypenko et al., 2021). Classifiers based on hand-crafted linguistic features rely upon tools such as parsers, and deep-learning-based classifiers might learn features that are specific to the training data but irrelevant to translationese (Amponsah-Kaakyire et al., 2022; Borah et al., 2023). These issues might hinder the generalization of the translationese detection.
Against this backdrop, we propose translationese index (T-index), the first method for graded and generalizable measurement of translationese, based on likelihood ratios of two contrastively fine-tuned LLMs, a low-translationese model and a high-translationese model. The method contrasts the distances between the sample and the training distribution of each model, estimating the level of translationese.
@@
We have several major findings: (1) Our proposed T-index is supported by/highly correlated with both pointwise and pairwise evaluations of translationese by human annotators. (2) The proposed method is generalizable to multiple genres and authors, even when the backbone LMs are trained on a single domain of only 1k samples, showing robustness and efficiency. (3) T-index has very low correlation with existing MT QE metrics, suggesting that it is a novel aspect not yet covered by existing metrics, and offers a new perspective for MT QE.
@@
The paper is organized as follows. We first introduce T-index in §2. Next, we describe the synthetic benchmark used for translationese measurement in §3. In §4, we evaluate the cross-domain generalization of T-index and several baselines on the synthetic benchmark. Finally, we evaluate T-index against human annotations for MTs in the wild in §5.
@@ 2 T-index: using likelihood ratios to measure translationese
Unlike previous research which was framed as a binary text-classification task between translated and non-translated text, we attempt to provide a continuous measure of translationese. For a source text $x$ and its translation $y$, we want to estimate the degree of translationese in $y$. The goal is to find a proper scoring function $f(x,y,\theta)$, parameterized by $\theta$. For each sample $(x,y)$, $f$ will yield a score that positively correlates with the degree of translationese.
As reported in the classification between original and translated texts, the supervised classifiers might learn confounding factors that are not necessarily related to translationese, such as the topic of the texts (Amponsah-Kaakyire et al., 2022). Such features might hinder the cross-domain generalization. Taking this into account, we set two confounding variables: (1) genre and (2) author, which means that the test samples might be sourced from different genres or translated by different authors. The automated translationese measurement is expected to tease out the translationese features from the stylistic variance brought by the genre and author.
@@
We show the validity of our proposed T-index in two steps. In the first step, we formulate the problem as classification, but instead of having translated and non-translated texts as the classes, we generate two classes of texts with extreme degrees of translationese, using carefully controlled prompts: one with very low-translationese, another with very high-translationese, and verify that human annotators agree with the low- vs. high-translationese distinction (§3). We then use T-index to perform the text classification to examine the discriminative power and generalizability of T-index between two extreme cases of translationese. (§4). The second step uses real-world translations and abandons the text-classification paradigm (§5). We first ask human annotators to rate the degree of translationese in the sampled translations. We then show that T-index is in line with human ratings using various methods.
@@ 3.2 Dataset statistics
We compare the statistics of previously studied linguistic features between the low and high-translationese samples, as shown in Table 1. These features are reported to be different between the original and translated texts in Chinese. We expect similar differences to be observed in the synthetic datasets.
@@
We conduct independent-sample t-tests to test the significance of the difference (see Table 1). Out of the 6 features, type-token ratio, function words, pronouns, and punctuations are in alignment with what has been reported in previous literature (Xiao and Hu, 2015; Hu et al., 2018; Hu and Kübler, 2021), suggesting that divergent linguistic patterns between our synthetic low- and high-translationese share similarity with that between the original and translated texts as previously reported and can be used as the starting point for studying T-index.
@@ 4 Classifying synthetic low- and high-translationese
Author (LLM as translator) Qwen2.5-72B-Instruct LLama3.3-70B-Instruct
Method / Genre OT (i.d.) Novel (o.o.d.) COCA (o.o.d.) OT (o.o.d.) Novel (o.o.d.) COCA (o.o.d.)
@@
DPO (Log-likelihood) $73.50_{~{}78.89}$ $76.44_{~{}82.76}$ $76.75_{~{}82.67}$ $75.00_{~{}80.70}$ $74.50_{~{}81.09}$ $75.88_{~{}82.90}$
Bradley-Terry RM $93.50_{~{}97.28}$ $88.41_{~{}93.97}$ $80.00_{~{}87.37}$ $88.67_{~{}94.53}$ $84.25_{~{}92.35}$ $83.83_{~{}91.56}$
XLM-RoBERTa $92.83$ $88.33$ $76.83$ $79.17$ $78.33$ $64.73$
SVM w/ ling. feats. 71.00 73.75 65.12 72.50 69.75 63.88
@@
Log-likelihood $83.33_{~{}89.56}$ $83.25_{~{}90.15}$ $79.83_{~{}86.61}$ $84.17_{~{}89.22}$ $80.83_{~{}88.12}$ $77.38_{~{}85.01}$
Entropy $71.17_{~{}76.04}$ $66.50_{~{}72.68}$ $66.54_{~{}71.14}$ $71.33_{~{}75.79}$ $66.25_{~{}70.77}$ $71.33_{~{}77.89}$
Fast-DetectGPT $80.33_{~{}87.26}$ $77.42_{~{}84.94}$ $78.33_{~{}85.41}$ $76.17_{~{}83.58}$ $75.08_{~{}81.01}$ $74.08_{~{}81.01}$
Maha. Distance $55.50_{~{}53.80}$ $53.08_{~{}51.76}$ $50.38_{~{}47.12}$ $56.17_{~{}54.31}$ $52.42_{~{}50.17}$ $50.62_{~{}47.69}$
Relative Maha. Dist. $77.17_{~{}84.13}$ $75.00_{~{}82.03}$ $75.38_{~{}81.28}$ $70.83_{~{}75.55}$ $70.83_{~{}76.97}$ $70.21_{~{}76.63}$
Trajectory Volatility $65.67_{~{}69.04}$ $59.42_{~{}59.19}$ $51.12_{~{}47.98}$ $57.67_{~{}57.58}$ $56.67_{~{}55.86}$ $52.79_{~{}51.40}$
@@
Log-likelihood $50.00_{~{}39.00}$ $50.50_{~{}33.12}$ $50.00_{~{}31.45}$ $50.00_{~{}32.89}$ $50.00_{~{}29.61}$ $50.00_{~{}28.01}$
Entropy $50.33_{~{}31.89}$ $50.00_{~{}33.13}$ $50.00_{~{}31.09}$ $50.00_{~{}30.07}$ $50.00_{~{}33.87}$ $50.00_{~{}26.83}$
Fast-DetectGPT $55.00_{~{}52.71}$ $50.42_{~{}45.60}$ $50.00_{~{}44.45}$ $51.00_{~{}45.98}$ $50.33_{~{}38.33}$ $50.17_{~{}39.50}$
Maha. Distance $55.33_{~{}54.51}$ $54.75_{~{}53.91}$ $54.67_{~{}55.19}$ $54.17_{~{}50.78}$ $54.83_{~{}53.67}$ $53.63_{~{}53.80}$
Relative Maha. Dist. $74.83_{~{}80.20}$ $73.50_{~{}79.08}$ $69.47_{~{}75.68}$ $69.33_{~{}72.76}$ $68.17_{~{}72.26}$ $65.37_{~{}69.39}$
Trajectory Volatility $58.00_{~{}58.60}$ $57.67_{~{}58.96}$ $61.20_{~{}64.00}$ $57.17_{~{}57.83}$ $59.25_{~{}60.84}$ $59.27_{~{}61.63}$
@@
Translationese Index $98.00_{~{}99.62}$ $95.17_{~{}98.99}$ $95.58_{~{}99.26}$ $92.67_{~{}96.43}$ $90.58_{~{}96.67}$ $89.63_{~{}95.94}$
Table 2: Results (accuracy reported with auroc as subscript) for the binary classification between low-translationese and high-translationese. All scoring and classifying models are trained on source-target pairs of Oliver Twist translated by Qwen2.5-72B-Instruct. i.d. indicates the in-domain test set, and all other columns denote cross-domain test sets (all results are the average across 3 random seeds).
We start with a binary classification task, distinguishing low-translationese translations from high-translationese translations. Specifically, we evaluate T-index along with several unsupervised (§4.1) and supervised (§4.2) baselines in the cross-domain settings.
@@ 4.1 Unsupervised baselines
To evaluate cross-domain generalization, we fix the sources sampled from Oliver Twist and translations generated by Qwen2.5-72B-Instruct as the only training domain. On paired samples from this domain, we SFT Qwen2.5-0.5B with the translation task to obtain two scoring models. One is fine-tuned on low-translationese translations, and the other is fine-tuned on high-translationese translations with the same set of hyperparameters. The scoring models are fine-tuned with a global batch size of 32 for 800 steps at a learning rate of 1e-06. Fine-tuning only takes around five minutes on four 3090 GPUs.
@@ 4.2 Supervised baselines
Aforementioned unsupervised methods do not rely on the labels of the training samples. We also include supervised baselines that are trained explicitly with labeled samples, including DPO-aligned Reward (Rafailov et al., 2023), Bradley-Terry Reward Model (Bradley and Terry, 1952; Ouyang et al., 2022), XLM-RoBERTa (Conneau et al., 2020), and SVM with linguistic features (Hu and Kübler, 2021).

### v2

@@ Abstract
Translationese refers to linguistic properties that usually occur in translated texts. Previous works study translationese by framing it as a binary classification between original texts and translated texts. In this paper, we argue that translationese should be graded instead of binary and propose the first measure for translationese—the translationese-index (T-index), computed from the likelihood ratios of two contrastively fine-tuned language models (LMs). We use synthesized translations and translations in the wild to evaluate T-index’s generalizability in cross-domain settings and its validity against human judgments. Our results show that T-index can generalize to unseen genres, authors, and language pairs. Moreover, T-index computed using two 0.5B LMs fine-tuned on only 1-5k pairs of synthetic data can effectively capture translationese, as demonstrated by alignment with human pointwise ratings and pairwise judgments. Additionally, the correlation between T-index and existing machine translation (MT) quality estimation (QE) metrics such as BLEU and COMET is low, suggesting that T-index is not covered by these metrics and can serve as a complementary metric in MT QE.
@@
Yikang Liu1thanks: Work done when Yikang Liu was an intern at Tongyi Lab. #Rui Wang and Hai Hu are corresponding authors., Wanyang Zhang2, Yiming Wang1, Jialong Tang3, Pei Zhang3, Baosong Yang3, Fei Huang3, Rui Wang1#, Hai Hu4# 1Shanghai Jiao Tong University   2Peking University 3Tongyi Lab   4City University of Hong Kong Correspondence: yikangliu@sjtu.edu.cn; wangrui12@sjtu.edu.cn; hu.hai@cityu.edu.hk
@@ 1 Introduction
Translationese refers to linguistic properties that are often introduced in the translation process that are different from those of texts originally written in that language (Gellerstam, 1986). While such properties are not inherently undesirable, they often lead to unnatural and non-native-like language that differs from idiomatic and authentic texts.
Translationese in translations, particularly machine translations (MTs), presents significant challenges in the era of large language models (LLMs). Many multilingual resources synthesized through MT have been reported as low quality (Kreutzer et al., 2022). Models trained on these “noisy” MT datasets often struggle to generalize effectively in real-world tasks that do not involve translation (Church et al., 2025). This issue extends even to high-resource languages: for Chinese, the second most resource-rich language, LLMs frequently produce unnatural texts resembling translationese in monolingual natural language generation tasks (Guo et al., 2024).
We believe this problem can be alleviated through a quantitative measure of translationese that enables selection of the most authentic and natural translation from multiple MT outputs. However, previous attempts to detect or measure translationese have failed to meet this goal. Existing approaches either develop binary classifiers to distinguish between translated and original texts (Baroni and Bernardini, 2005; Koppel and Ordan, 2011; Volansky et al., 2015; Hu and Kübler, 2021; Pylypenko et al., 2021), or rely on distributional statistics computed across batches of texts rather than individual samples (Freitag et al., 2022; Guo et al., 2024; Li et al., 2025; Flamich et al., 2025), in which original texts are regarded as no-translationese distribution. The former approach lacks continuous measurement capabilities, while the latter cannot score individual translations.
In this paper, we want to find a graded and generalizable measurement of translationese. To this end, we first reframe how translationese should be discovered. We argue that the degree of translationese should be directly compared among translations. This new framework offers two key benefits: first, graded comparisons can be more easily observed, and second, it isolates confounding factors that arise in classification between original and translated texts, thereby enabling better generalization.
@@
Under the new problem formulation, we propose translationse index (T-index) and compare it with several supervised and unsupervised baselines on synthetic and human-annotated data. We have several major findings: (1) The proposed method is generalizable to multiple genres, authors, and language pairs, even when the backbone LMs are fine-tuned with only 1k synthesized samples. (1) T-index is correlated with both pointwise (Pearson’s $r=0.418$) and pairwise evaluations of translationese by expert human annotators. (3) T-index has very low correlation with existing MT quality estimation (QE) metrics, suggesting that it might be a novel aspect not yet covered by existing metrics.
@@
The paper is organized as follows. We first introduce our new problem formulation and T-index in §2. Next, we describe the multi-genre synthetic benchmark used for translationese measurement in §3. In §4, we evaluate the cross-domain generalization of T-index and several baselines on the synthetic benchmark. Finally, we evaluate T-index against human annotations for MTs in the wild in §5.
@@ 2 T-index: using likelihood ratios to measure translationese
Unlike previous research, which was framed as a binary text-classification task between translated and non-translated text, we compare different translations of the same source text, attempting to provide a continuous measure of translationese. The new framing is motivated by the following rationales:
  1. 1.
    Whether the texts are translated or not does not determine the degree of translationese, i.e., translated texts can also be authentic, and non-translated ones might be unnatural.
  2. 2.
    The binary-classification between translated and non-translated texts is sometimes confounded by features unrelated to translationese, such as the topic of the texts (Amponsah-Kaakyire et al., 2022; Borah et al., 2023), hindering the discovery of translationese-specific features.
In the new formulation, we do not consider non-translated texts as having “no translationese”, but instead aim to measure the degree of translationese of each individual translation.
Specifically, for a source text $x$ and its translation $y$, we want to estimate the degree of translationese in $y$. The goal is to find a proper scoring function $f(x,y,\theta)$, parameterized by $\theta$. For each sample $(x,y)$, $f$ will yield a score that directly predicts the degree of translationese.
In order to make the scoring function genuinely generalizable, we consider two possible confounding variables: (1) genre and (2) author. Ideally, our scoring function should only capture abstract translationese-specific features, rather than genre- or author-related textual features of the translations. Thus, to test whether the scoring function is generalizable and robust, we include test samples from different genres or translated by different authors.
@@
We show the validity of T-index in two steps. In the first step, we formulate the problem as classification, but instead of having translated and non-translated texts as the classes, we generate two classes of texts with extreme degrees of translationese, using carefully controlled prompts: one with very low translationese, the other with very high translationese, and verify that human annotators agree with the low- vs. high-translationese distinction (§3). We then use T-index to perform text classification to examine the discriminative power and generalizability of T-index for texts on the two ends of translationese (§4). The second step uses real-world translations and abandons the text-classification paradigm (§5). We first ask human annotators to rate the degree of translationese in the sampled translations. We then show that T-index is in line with human ratings using various methods.
@@ 3.2 Dataset statistics
We compare the statistics of previously studied linguistic features between the low and high-translationese samples, as shown in Table 1. These features are reported to differ between original and translated texts in Chinese. We expect similar differences in our synthetic dataset.
@@
We conduct independent-sample t-tests to test the significance of the difference (see Table 1). Out of the 6 features, type-token ratio, function words, pronouns, and punctuations are in alignment with what has been reported in previous literature on Chinese (Xiao and Hu, 2015; Hu et al., 2018; Hu and Kübler, 2021), suggesting that divergent linguistic patterns between our synthetic low- and high-translationese share similarity with that between the original and translated texts as previously reported and can be used as the starting point for studying T-index.
@@ 3.3 Human annotation
3.4 More language pairs
We also synthesize French and Germany translations on the basis of the same English source texts with Qwen2.5-72B-Instruct, repeating the generation scheme used for English-Chinese pairs, resulting in 14 en-de and 14 en-fr test sets. Note that we have not validated the quality of generated samples as carefully as what we did to en-zh pairs, and can only provide a preliminary exploration of the cross-linguistic generalization of various measures of translationese with the two language pairs.
@@ 4 Classifying synthetic low- and high-translationese
Language pair en-zhid en-deood en-frood
Author (LLM as translator) Qwen2.5-72B-INSTid LLama3.3-70B-INSTood Qwen2.5-72Bid
Method $\backslash$ Genre OTid Novelood COCAood OTid Novelood COCAood All
@@
DPO (Log-likelihood) $89.2_{~95.8}$ $86.9_{~94.5}$ $82.3_{~89.5}$ $89.2_{~95.5}$ $86.8_{~94.4}$ $80.5_{~88.2}$ $77.9_{~85.6}$ $82.6_{~89.7}$
Bradley-Terry RM $94.7_{~98.4}$ $87.6_{~94.2}$ $87.0_{~93.9}$ $87.3_{~94.3}$ $89.9_{~95.6}$ $87.2_{~93.8}$ $72.9_{~79.8}$ $76.5_{~84.6}$
XLM-RoBERTa $95.0_{~\text{---}}$ $83.6_{~\text{---}}$ $82.1_{~\text{---}}$ $81.8_{~\text{---}}$ $87.9_{~\text{---}}$ $68.9_{~\text{---}}$ $56.8_{~\text{---}}$ $60.7_{~\text{---}}$
SVM w/ ling. feats. $71.0_{~\text{---}}$ $73.7_{~\text{---}}$ $65.1_{~\text{---}}$ $72.5_{~\text{---}}$ $69.7_{~\text{---}}$ $63.8_{~\text{---}}$ $\text{---}_{~\text{---}}$ $\text{---}_{~\text{---}}$
@@
Log-likelihood $80.1_{~87.1}$ $79.2_{~85.9}$ $79.0_{~86.1}$ $80.1_{~87.0}$ $77.6_{~84.8}$ $76.9_{~84.4}$ $71.0_{~77.7}$ $75.3_{~81.3}$
Entropy $73.8_{~77.0}$ $64.8_{~71.0}$ $69.1_{~73.7}$ $71.6_{~74.4}$ $65.4_{~71.5}$ $71.0_{~77.9}$ $60.3_{~63.9}$ $61.1_{~65.4}$
Fast-DetectGPT $74.3_{~80.8}$ $72.1_{~78.9}$ $73.1_{~80.5}$ $72.0_{~79.2}$ $72.0_{~79.7}$ $68.5_{~74.9}$ $70.2_{~76.6}$ $76.0_{~83.6}$
Maha. Distance $55.1_{~54.3}$ $52.4_{~51.0}$ $50.2_{~47.2}$ $54.8_{~53.6}$ $52.7_{~51.3}$ $50.3_{~47.5}$ $51.6_{~50.8}$ $51.0_{~47.7}$
Relative Maha. Dist. $78.3_{~86.0}$ $70.8_{~77.9}$ $76.0_{~83.3}$ $72.8_{~76.9}$ $73.1_{~80.4}$ $70.8_{~77.4}$ $50.1_{~47.9}$ $50.5_{~49.5}$
Trajectory Volatility $63.6_{~69.9}$ $58.5_{~59.2}$ $52.1_{~49.6}$ $60.1_{~60.0}$ $60.1_{~62.0}$ $53.1_{~52.0}$ $51.1_{~49.6}$ $50.7_{~47.9}$
@@
Log-likelihood $50.0_{~31.5}$ $50.0_{~28.6}$ $50.0_{~25.6}$ $50.0_{~26.9}$ $50.1_{~31.0}$ $50.0_{~24.4}$ $50.0_{~27.4}$ $50.0_{~24.9}$
Entropy $50.3_{~29.5}$ $50.0_{~32.8}$ $50.0_{~28.8}$ $50.0_{~29.8}$ $50.1_{~32.6}$ $50.0_{~25.6}$ $50.0_{~38.3}$ $50.0_{~37.6}$
Fast-DetectGPT $53.5_{~49.0}$ $50.1_{~39.3}$ $50.0_{~42.2}$ $51.1_{~40.9}$ $50.0_{~41.9}$ $50.0_{~40.0}$ $50.0_{~28.8}$ $50.1_{~22.2}$
Maha. Distance $56.3_{~55.0}$ $4.6_{~54.3} 55$$1_{~56.1} 55.$$8_{~51.5} 55.$$5.8_{~55.2}$ $54.0_{~54.4}$ $51.3_{~49.6}$ $53.3_{~53.2}$
Relative Maha. Dist. $74.8_{~81.9}$ $69.1_{~74.9}$ $72.2_{~78.5}$ $71.3_{~74.5}$ $70.0_{~77.0}$ $66.6_{~72.3}$ $50.2_{~48.9}$ $50.7_{~49.9}$
Trajectory Volatility $58.8_{~59.6}$ $57.3_{~58.2}$ $61.8_{~65.3}$ $0_{~56.6} 57.75$$7_{~58.7}$ $60.0_{~62.4}$ $51.8_{~50.8}$ $51.7_{~50.8}$
@@
Translationese Index $95.8_{~99.2}$ $92.7_{~97.8}$ $95.2_{~98.8}$ $93.0_{~97.9}$ $93.5_{~98.3}$ $90.5_{~96.2}$ $79.0_{~85.5}$ $82.3_{~90.0}$
Table 2: Results (accuracy reported with auroc as subscript) for the binary classification between low-translationese and high-translationese. All scoring and classifying models are trained on source-target pairs of Oliver Twist translated by Qwen2.5-72B-Instruct. id indicates the in-domain test set, and all other columns denote cross-domain test sets (all results are the average across 3 random seeds).
We start with a binary classification task, distinguishing translations of low-translationese from those with high-translationese. Specifically, we evaluate T-index along with several unsupervised (§4.1) and supervised (§4.2) baselines in the cross-domain settings.
@@ 4.1 Unsupervised baselines
On paired samples from the training domain, we SFT Qwen2.5-0.5B with the translation task to obtain two scoring models. One is fine-tuned on low-translationese translations, and the other on high-translationese translations.
@@ 4.2 Supervised baselines
The methods above do not rely on the labels of the training samples. We also include supervised baselines trained explicitly with labeled samples, including DPO-aligned Reward (Rafailov et al., 2023, 2024), Bradley-Terry Reward Model (Bradley and Terry, 1952; Ouyang et al., 2022), XLM-RoBERTa (Conneau et al., 2020), and SVM with linguistic features (Hu and Kübler, 2021).